OpenStudy (anonymous):

Which of the following is the solution to the equation 25^(z + 2) = 125

OpenStudy (anonymous):

\[(z+2)\ln 25 = \ln 125 \Longrightarrow z = \frac{\ln 125}{\ln 25} - 2 = \frac{3\ln 5}{\ln 5} = 3-2 = 1\]

OpenStudy (anonymous):

the choices were z=-2.5 z=5.5 z=3.5 z=-0.5

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):